Solve the partial differential equation :
z (z^2-xy )(px-qy) = x^4
1
Expert's answer
2020-10-06T18:53:01-0400

Given Equation is "z(z^2-xy) (px-qy) = x^4"


then "px-qy = \frac{x^4}{z(z^2-xy)}"


comparing with general equation "Pp+Qq=R"


Then "\frac{dx}{x} = \frac{dy}{-y} = \frac{dz}{ \frac{x^4}{z(z^2-xy)}}"


Taking first two terms,

"\frac{dx}{x} = \frac{dy}{-y}"

Integrating both sides,

"\int \frac{dx}{x} = \int\frac{dy}{-y}"

"log (x) = -log(y) + logC"

"xy = C" (1)



Taking first and last term,

"\frac{dx}{x} = \frac{dz}{ \frac{x^4}{z(z^2-C)}}"


Integrating both sides, "\int x^3{dx} =\int z(z^2-C){dz}"


"x^4 = z^4-2z^2C + C' = z^4-2xyz^2 + C'"

"x^4 -z^4+2xyz^2=C'" (2)



So required solution is

"f(xy,x^4-z^4+2xyz^2) = 0"